# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v3

@@ 5.4 Protocol Profile Comparison
Table 6 compares the three deployment profiles with experimental measurements for ML-DSA-65, $(n=5,t=3)$ using our optimized Rust implementation.
@@
Table 6: Protocol profile comparison (ML-DSA-65, $n=5$, $t=3$, Rust implementation). P3+ = 1 signer step + 1 server round; signers precompute nonces offline.
@@ 5.5 Comparison with Prior Work
Table 7: Comparison with prior work. Times for prior schemes are from their original papers; our times are from our Rust implementation (ML-DSA-65, $n=5$, $t=3$). Our $\approx$21–45% success rate is comparable to ML-DSA’s expected $\approx$20–25%. P3+ = 1 signer step + 1 server round.
@@ 5.8 Preprocessing Analysis
Profile P2 (MPC) benefits significantly from preprocessing. Table 9 shows the amortization effect.
@@ 5.12 FIPS 204 Compliance Verification
We verified our implementation against FIPS 204 specifications. Parameter verification confirmed all 19 FIPS 204 constants ($q=8380417$, $n=256$, $d=13$, etc.). Algorithm correctness testing covered KeyGen, Sign, Verify, and all auxiliary functions (HighBits, LowBits, MakeHint, UseHint, NTT, sampling) with 117 test cases. Bound verification confirmed all signatures satisfy $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ and hint weight $\leq\omega$.
@@ 5.13 Code Availability
We provide two implementations:
  • Python reference ($\approx$3500 lines): Readable implementation for protocol understanding and verification.
[…]
@@
  • Comprehensive protocol test suite with 40+ integration tests (separate from 117 FIPS compliance tests)
@@ 6.1 Distributed Key Generation
Our main protocol assumes a trusted dealer. A standard DKG protocol [GJKR07] with Feldman commitments [Feldman87] can eliminate this assumption: each party contributes a random sharing of local secrets, and parties aggregate their shares. For lattice-specific VSS/DKG, see [FS24]. The final secret $\mathbf{s}_{1}=\sum_{i}\mathbf{s}_{1}^{(i)}$ is unknown to any individual party. See Supplementary Material for the full protocol.
@@ 6.2 Optimistic Blame Protocol
    TEE requests each signer to reveal nonce seeds and mask inputs (sent only to TEE)
@@
    TEE recomputes $\mathbf{y}_{i}$, $\mathbf{w}_{i}$, masks, and verifies consistency with submitted values
@@ 6.4 Security Analysis of Profile P2
for ML-DSA-65 parameters.
@@ 6.5.1 UC Security Analysis of Profile P3+
  • Input: Receive $(\mathbf{w},\mathbf{S}_{1})$ from $\mathsf{CP}_{1}$ and $\mathbf{S}_{2}$ from $\mathsf{CP}_{2}$
  • Compute: $\mathbf{w}^{\prime}=\mathbf{w}-\mathbf{S}_{1}-\mathbf{S}_{2}\bmod q$;  $\mathsf{result}=\bigwedge_{j}(|\mathbf{w}^{\prime}_{j}|_{q}<\gamma_{2}-\beta)$
  • Output: Send $\mathsf{result}$ to both parties
@@ Conclusion.
    Privacy: Masks hide individual shares; 2PC reveals only result bit
@@ 6.7 Additional Extensions
Adaptor Signatures.
For atomic swaps: parties compute $\mathbf{w}^{\prime}=\mathbf{w}+\mathbf{R}$ with adaptor point $\mathbf{R}$; pre-signature can be completed by anyone knowing witness $\mathbf{r}$ with $\mathbf{A}\mathbf{r}=\mathbf{R}$.
@@ 7 Conclusion
We have presented Masked Threshold ML-DSA, a threshold signature scheme for FIPS 204 ML-DSA that supports arbitrary thresholds while maintaining standard signature sizes and verification compatibility with existing FIPS 204 verifiers.
@@ 7.2 Summary of Contributions
Our main technical contribution is masked Lagrange reconstruction, which adapts pairwise-canceling masks [CGGMP20] to the lattice setting. This approach addresses three challenges absent in ECDSA: rejection sampling bounds, r0-check key leakage, and Irwin-Hall nonce security. The key insight is ensuring only the sum needs to be small, not individual contributions.
We instantiate this technique in three deployment profiles. Profile P1 (TEE-Assisted) achieves 3-round signing with EUF-CMA security under Module-SIS (Theorem 4.2). Profile P2 (Fully Distributed) uses MPC to achieve 5-round signing with dishonest-majority and UC security (Theorem 6.2). Profile P3+ (Semi-Async 2PC) achieves 2 logical rounds (22 ms for 3-of-5) with UC security under a 1-of-2 CP honest assumption (Theorem 6.5).
@@ 7.3 Limitations and Future Work
Each deployment profile involves different trust assumptions: Profile P1 requires TEE/HSM trust, Profile P2 eliminates hardware trust at the cost of 5 MPC rounds, and Profile P3+ requires 1-of-2 CP honesty with 2 logical rounds. Our protocol requires $|S|\geq T+1$ signers to ensure at least two honest parties for mask privacy, which is inherent to pairwise-mask schemes [CGGMP20]. The aggregated nonce follows an Irwin-Hall rather than uniform distribution; while the security loss is negligible ($R_{2}^{\epsilon}\leq 1.003$ for $|S|\leq 17$, $\leq 1.05$ for $|S|\leq 33$; see Appendix D), this is a deviation from single-signer ML-DSA. Finally, we assume a synchronous model where all parties are online simultaneously during signing.
@@
Several directions merit further investigation. Distribution-preserving variants that maintain uniform nonces would eliminate the Irwin-Hall deviation entirely. Two-round protocols would reduce latency, though at the cost of additional complexity. Efficient malicious security via lightweight zero-knowledge proofs could strengthen Profile P1. Finally, formal verification in EasyCrypt would provide machine-checked security guarantees.

### v4

@@ 5.4 Protocol Profile Comparison
Table 6 compares the three deployment profiles with experimental measurements for ML-DSA-65, (N=5,T=3)(N\!=\!5,T\!=\!3) using our optimized Rust implementation.
@@
Table 6: Protocol profile comparison (ML-DSA-65, N=5N\!=\!5, T=3T\!=\!3, Rust implementation). P3+ = 1 signer step + 1 server round; signers precompute nonces offline.
@@ 5.5 Comparison with Prior Work
Table 7: Comparison with prior work. Times for prior schemes are from their original papers; our times are from our Rust implementation (ML-DSA-65, N=5N\!=\!5, T=3T\!=\!3). Our $\approx$21–45% success rate is comparable to ML-DSA’s expected $\approx$20–25%. P3+ = 1 signer step + 1 server round.
@@ 5.8 Preprocessing Analysis
All profiles benefit from preprocessing. The Shamir nonce DKG (Section 3.4) adds one offline round of $O(|S|^{2})$ point-to-point messages ($\approx$3.8 KB each for ML-DSA-65) before each signing session. In practice, multiple nonce sharings can be precomputed during idle time and consumed on demand, amortizing the DKG cost across signing sessions.
Profile P2 (MPC) benefits from additional preprocessing. Table 9 shows the amortization effect.
@@
Deployment Latency Estimates.
Our benchmarks measure computation time on a single machine. In distributed deployments, end-to-end latency is approximately $t_{\mathsf{compute}}+\text{rounds}\times\text{RTT}$. On a LAN (1 ms RTT), P1 takes $\approx$10 ms and P3+ $\approx$19 ms for $(3,5)$-threshold. Over a WAN (20 ms RTT), P1 rises to $\approx$67 ms and P2 to $\approx$120 ms. Over intercontinental links (100 ms RTT), P2 reaches $\approx$520 ms, while P3+ remains under 220 ms due to its 2-round structure. Detailed estimates appear in the supplementary material. P3+’s reduced round count makes it the most network-friendly profile for geographically distributed deployments.
@@ 5.12 FIPS 204 Compliance Verification
We verified our implementation against FIPS 204 specifications. Parameter verification confirmed all FIPS 204 constants ($q=8380417$, $n=256$, $d=13$, $\omega=55$, etc.). Algorithm correctness testing covered KeyGen, Sign, Verify, and all auxiliary functions (HighBits, LowBits, MakeHint, UseHint, NTT, SampleInBall, Power2Round, w1Encode) with FIPS 204 Known-Answer Tests (KAT) verifying each building block against specification-derived test vectors. Bound verification confirmed all signatures satisfy $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ and hint weight $\leq\omega$.
@@ 5.13 Code Availability
Our primary implementation is in Rust ($\approx$5500 lines core + $\approx$4000 lines tests), with production-ready performance including parallel processing. An earlier Python prototype was used during development but is no longer maintained; the Rust implementation is the reference.
@@
  • Comprehensive test suite: FIPS 204 KAT tests for all building blocks plus protocol integration tests across all profiles
@@ 6.1 Distributed Key Generation
Our protocol employs two distinct DKG protocols. The key DKG (executed once during setup) eliminates the trusted dealer assumption: each party contributes a random sharing of local secrets sampled from the short-coefficient distribution $\chi_{\eta}$, parties aggregate their shares, and Feldman commitments [Feldman87] provide robustness against malicious participants. The final secret $\mathbf{s}_{1}=\sum_{i}\mathbf{s}_{1}^{(i)}$ is unknown to any individual party. For lattice-specific VSS/DKG, see [FS24, GJKR07].
The nonce DKG (executed before each signing session) is our primary technical contribution. As described in Section 3.4, each party generates a degree-(T1)(T\!-\!1) polynomial whose higher-degree coefficients are sampled uniformly from $R_{q}^{\ell}$, producing a fresh Shamir sharing of the signing nonce. This structural choice enables statistical share privacy (Theorem 4.4). The nonce DKG can be preprocessed offline and amortized across multiple signing sessions; see the Supplementary Material for the full protocol details.
@@ 6.2 Optimistic Blame Protocol
    TEE requests each signer to reveal their nonce DKG polynomials and mask inputs (sent only to TEE)
@@
    TEE recomputes nonce shares $\mathbf{y}_{i}$, commitments $\mathbf{w}_{i}$, responses $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$, and masks, verifying consistency with submitted values
@@ 5 Output
UC Security of 5-Round Optimization.
The 5-round protocol is obtained from the 8-round base protocol (Appendix H) via two standard MPC composition steps. First, combiner-mediated commit-then-open merges sequential commit and open rounds (Rounds 3–4 $\to$ Round 2): parties send $(\mathsf{Com}_{i},\mathsf{masked}_{i},\mathsf{rand}_{i})$ to a designated combiner who releases all openings simultaneously; the combiner is an ordinary party, and SPDZ MAC checks detect any tampering, so this is UC-equivalent to a sequential broadcast [CLOS02]. Second, parallel Beaver-triple-based comparison merges the sequential comparison rounds (Rounds 6–7 $\to$ Round 4): all 1536 coefficient comparisons execute in parallel (Layer 1), followed by a single-round AND-tree aggregation (Layer 2). Each optimization replaces a sub-protocol with a functionally equivalent variant; by the UC composition theorem [Canetti01], the composed 5-round protocol UC-realizes the same ideal functionality. The hybrid argument in Appendix H ($\mathsf{H}_{0}$$\mathsf{H}_{6}$) carries over with the round assignment remapped, since each hybrid step depends on the functionality of each sub-protocol (SPDZ, edaBits, binary circuits), not on its round structure. The total security loss remains $\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$.
@@ 6.4 Security Analysis of Profile P2
for ML-DSA-65 parameters, where $\epsilon_{\mathsf{SPDZ}}\leq O(m)/q<2^{-12}$ with $m=nk=1536$ coefficients and $\epsilon_{\mathsf{edaBits}}<2^{-12}$ with cut-and-choose parameter $B=64$.
@@ 6.5 Profile P3+: Semi-Async 2PC-Assisted Signing
Hint Computation.
The 2PC circuit is extended to output not only the r0-check result bit, but also the hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$ when the check passes. The hint computation requires $c\mathbf{s}_{2}$ (to derive $c\mathbf{t}_{0}$), which is available inside the 2PC as $\mathbf{S}_{1}+\mathbf{S}_{2}$. Since $\mathbf{h}$ is part of the public signature, outputting it does not compromise privacy. The circuit adds $O(nk)$ comparison gates for $\mathsf{MakeHint}$, negligible relative to the r0-check circuit.
Combiner and Challenge Integrity.
[…]
@@ 6.5.1 UC Security Analysis of Profile P3+
  • Public input: $(\mathbf{A},c,\mathbf{t}_{1},\mathbf{z})$ (known to both parties and the combiner)
  • Private input: Receive $(\mathbf{w},\mathbf{S}_{1})$ from $\mathsf{CP}_{1}$ and $\mathbf{S}_{2}$ from $\mathsf{CP}_{2}$
  • Compute: $\mathbf{w}^{\prime}=\mathbf{w}-\mathbf{S}_{1}-\mathbf{S}_{2}\bmod q$;  $\mathsf{result}=\bigwedge_{j}(|\mathbf{w}^{\prime}_{j}|_{q}<\gamma_{2}-\beta)$. If $\mathsf{result}=1$: derive $c\mathbf{t}_{0}=(\mathbf{A}\mathbf{z}-\mathbf{w})+(\mathbf{S}_{1}+\mathbf{S}_{2})-c\mathbf{t}_{1}\cdot 2^{d}$ and compute $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$
  • Output: Send $\mathsf{result}$ to both parties. If $\mathsf{result}=1$, additionally output $\mathbf{h}$
@@ Conclusion.
    Privacy: Response shares $\mathbf{z}_{i}$ enjoy statistical privacy via Shamir nonce DKG (Theorem 4.4); commitment and r0-check values hidden by pairwise masks; 2PC reveals only the result bit
@@ 7 Conclusion
We have presented Masked Threshold ML-DSA, the first threshold signature scheme for FIPS 204 ML-DSA that achieves statistical share privacy (no computational assumptions) with arbitrary thresholds while maintaining standard signature sizes and verification compatibility with existing FIPS 204 verifiers. In coordinator-based profiles (P1, P3+), signing sets of size $|S|=T$ suffice; the fully distributed profile (P2) retains $|S\setminus C|\geq 2$ for commitment privacy.
@@ 7.2 Summary of Contributions
Our primary technical contribution is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, producing a degree-(T1)(T\!-\!1) Shamir sharing that structurally matches the long-term secret. Because the adversary observes only T1T-1 evaluations of each honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s response $\mathbf{z}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$, requiring no computational assumptions (Theorem 4.4). The bounded-nonce constraint prevents perfect uniformity (Remark 4.6), but the min-entropy guarantee is more than sufficient for practical security. In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for response privacy: signing sets of size $|S|\geq T$ suffice, matching the information-theoretic minimum for Shamir reconstruction. Individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator (not broadcast), preventing a key-recovery attack via the injective matrix $\mathbf{A}$ (Remark 4.1). Commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.9), which applies to P2’s broadcast model.
As a secondary technique, we introduce pairwise-canceling masks adapted from ECDSA [CGGMP20] to the lattice setting, used for commitment values $\mathbf{W}_{i}$ and r0-check shares $\mathbf{V}_{i}$. This addresses three challenges absent in ECDSA: rejection sampling bounds, r0-check key leakage, and Irwin-Hall nonce security.
We instantiate these techniques in three deployment profiles with complete UC proofs. Profile P1 (TEE-Assisted) achieves 3 online rounds plus 1 offline preprocessing round with EUF-CMA security under Module-SIS (Theorem 4.2). Profile P2 (Fully Distributed) uses MPC to achieve 5-round signing with dishonest-majority and UC security (Theorem 6.2). Profile P3+ (Semi-Async 2PC) achieves 2 logical rounds (22 ms for 3-of-5) with UC security under a 1-of-2 CP honest assumption (Theorem 6.5).
@@ 7.3 Limitations and Future Work
Each deployment profile involves different trust assumptions: Profile P1 requires TEE/HSM trust, Profile P2 eliminates hardware trust at the cost of 5 MPC rounds, and Profile P3+ requires 1-of-2 CP honesty with 2 logical rounds. The nonce DKG adds one offline preprocessing round of $O(N^{2})$ point-to-point messages ($\approx$3.8 KB each) before each signing session; while this can be amortized by precomputing nonce batches during idle time, it does increase protocol complexity relative to purely additive nonce sharing. The aggregated nonce follows an Irwin-Hall rather than uniform distribution; while the per-coordinate Rényi divergence is small ($R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$), the full-vector composition $(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ grows to $\approx 2^{5.4}$ for $|S|=17$ ($\approx$3-bit loss) and $\approx 2^{75}$ for $|S|=33$ ($\approx$38-bit loss); see Theorem 4.11. For the pairwise masks used on commitment and r0-check values, $|S\setminus C|\geq 2$ honest parties are needed. In coordinator-based profiles (P1, P3+), masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast, so $|S\setminus C|=1$ suffices for response privacy; in P2’s broadcast model, the existing mask-hiding condition $|S\setminus C|\geq 2$ applies to all values. Finally, we assume a synchronous model where all parties are online simultaneously during signing (though P3+ relaxes this to semi-asynchronous participation).
@@
Several aspects of our construction invite further investigation, ranging from strengthening the security model to broadening the parameter space and exploring applied deployment scenarios.
The most natural theoretical extension is adaptive security. Our analysis assumes a static adversary that selects which parties to corrupt before protocol execution begins; extending the results to the adaptive setting—where the adversary may corrupt parties during the protocol based on observed transcripts—remains open. The Shamir nonce DKG introduces a subtle obstacle: each honest party’s polynomial $\mathbf{f}_{h}$ has its leading coefficient $\mathbf{a}_{h,T-1}$ sampled uniformly, but the remaining coefficients are constrained by the constant term $\hat{\mathbf{y}}_{h}$ and the evaluation points. Under adaptive corruption, the simulator must equivocate these polynomials after the fact, which may require an erasure model (parties delete intermediate state after each round) or a non-committing encryption wrapper around the DKG messages. The analogous question for ECDSA threshold signatures was resolved by Canetti et al. [CGGMP20] using UC-secure commitments; whether a similar approach suffices in the lattice setting, where the polynomial structure imposes additional algebraic constraints, is unclear.
A related direction concerns the strength of the abort guarantee. All three profiles currently achieve security-with-abort: a malicious party can force the protocol to output $\bot$, but cannot produce an invalid signature. Profile P2’s SPDZ-based MPC naturally supports identifiable abort—when a MAC check fails, the cheating party can be pinpointed and excluded—but we have not formalized this property. A stronger goal is guaranteed output delivery, where the protocol always produces a valid signature provided sufficiently many honest parties participate. This is achievable in the honest-majority setting via broadcast protocols [CLOS02], and Profile P2’s framework can be augmented with a fair reconstruction phase when $|S|\geq 2T-1$. Such an upgrade would be particularly relevant for custody and government applications where signing requests must not be silently suppressed.
Our appendix sketches a share refresh mechanism based on adding shares of zero, but a complete treatment of proactive threshold ML-DSA requires addressing several interacting concerns: maintaining signing availability during the refresh epoch, ensuring consistency at epoch boundaries (parties must agree on which shares are current), and formally modeling a mobile adversary that corrupts different subsets in different epochs subject to a per-epoch threshold. The Shamir structure of our key shares is well-suited to proactive protocols—refresh amounts to adding a fresh degree-(T1)(T\!-\!1) zero-sharing—but the interaction between nonce DKG freshness and proactive share updates has not been analyzed. In particular, if a party’s long-term share $\mathbf{s}_{1,i}$ is refreshed between the nonce DKG and the signing round, the Lagrange reconstruction $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}$ may fail unless both the nonce and secret shares are refreshed consistently within the same epoch.
On the protocol side, three open problems stand out. The most practically significant concerns eliminating the Irwin-Hall security gap for large signing sets. Our proven security bound degrades with $|S|$ because the aggregated nonce $\mathbf{y}=\sum_{i}\lambda_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than the uniform distribution assumed by FIPS 204: the full-vector Rényi divergence $(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ grows from $\approx 2^{5.4}$ ($|S|=17$, $\approx$3-bit security loss) to $\approx 2^{75}$ ($|S|=33$, $\approx$38-bit loss). For practical deployments with $|S|\leq 17$—covering the vast majority of HSM clusters, custody quorums, and certificate authority configurations—the proven bound provides full 128-bit security. For larger committees, the distribution-preserving rejection sampling technique of Raccoon [Raccoon2024] offers a principled solution: after computing the aggregated nonce, parties collectively evaluate a rejection function that flattens the Irwin-Hall distribution to uniform, yielding $R_{2}^{\mathsf{vec}}=1$ (no security loss) at the cost of an additional interactive rejection round and a reduced per-attempt success rate. The key technical challenge in the threshold setting is that $\mathbf{y}$ is secret-shared—no single party knows the full nonce—so the rejection decision must be computed via MPC over the shares, adding communication complexity proportional to the rejection function’s circuit size. Whether a lightweight, non-interactive variant exists (e.g., each party independently rejecting based on local share statistics) remains open.
Second, this work focuses exclusively on ML-DSA-65 (NIST Security Level 3). The other two standardized parameter sets present distinct challenges: ML-DSA-44 (Level 2) has a smaller masking range $\gamma_{1}=2^{17}$, which tightens the per-coordinate Rényi bound by a factor of $\approx 16$, potentially pushing $R_{2}^{\epsilon,\mathsf{coord}}$ above $1.05$ for moderate signing sets and making the full-vector composition correspondingly more severe. Conversely, ML-DSA-87 (Level 5) uses $k=8,\ell=7$, increasing the r0-check circuit from 1536 to 2048 coefficients and the nonce DKG message size from 3.8 KB to 7.2 KB. A systematic parameter study with benchmarks across all three levels would be valuable for deployment planning.
Third, in P2’s broadcast model, the pairwise-canceling masks on $\mathbf{W}_{i}$ and $\mathbf{V}_{i}$ require $|S\setminus C|\geq 2$ for computational hiding (Lemma 4.9). In the extreme case $T=N$ (unanimous signing), a single honest party’s commitment and r0-check values are not hidden in P2. (In P1 and P3+, the coordinator-only communication model protects $\mathbf{W}_{i}$ even with $|S\setminus C|=1$; see Remark 4.1.) Replacing pairwise PRF masks with a non-interactive zero-sharing scheme—where each party’s mask is derived from a common reference string rather than pairwise seeds—or employing a verifiable PRF construction could close this gap, yielding uniform $|S|\geq T$ privacy across all revealed values. Additionally, achieving true information-theoretic privacy (SD =0=0) for the response shares with $|S|=T$ remains an open problem: the bounded-nonce constraint fundamentally prevents $\mathbb{Z}_{q}$-uniformity (Remark 4.6), and it is unclear whether alternative nonce generation techniques can bypass this barrier without requiring $|S|\geq T+1$.
Finally, on the applied side, threshold ML-DSA is a natural building block for post-quantum decentralized identity systems. The W3C Decentralized Identifier specification [W3CDID] requires controller key rotation and recovery mechanisms that interact non-trivially with threshold key management: a DID controller update must atomically rotate the threshold public key while preserving the mapping between DID documents and verification methods. Combining our protocol with a threshold-friendly DID resolution layer and a key recovery flow based on social recovery via Shamir shares would yield an end-to-end post-quantum identity system suitable for regulated environments.